Find the appropriate height of the building of a wire is connected to the top of the building is 450 feet and the distance from the building to the wire is 200 feet

Answer by Boreal(15235)

The 450 feet is the hypotenuse, 200 feet is one of the legs, and the height of the building is the other leg.
200^2+x^2=450^2
40000+x^2=202500
x^2=162500
x=403.11 feet